OpenStudy (anonymous):

Which line shows the first error in the solution (1) 5(x-5)=15 (2) 5x- 25 15 (3) 5x - 25 = 15 - 25 (4) 5x = -10 (5) 5x/5 = -10/5 (6) X= -2 A. Line 4 B line 5 C line 2 D line 3

OpenStudy (irishboy123):

well you're missing an equals sign in (2) i suspect that that's just a typo but if you need to get that level of understanding right before proceeding....

OpenStudy (anonymous):

it would be C.

OpenStudy (anonymous):

Line 2 because you have to subtrct 5 from both sides

OpenStudy (anonymous):

I would say C as well

OpenStudy (lightn1ng):

It's D.

OpenStudy (anonymous):

D If line 2's missing equals are just a typo.For the 25 to be anulled ,to isolate the 25,one should null the subtraction,then adding its valor(of course you could still take 25 if you want,then it needs to stay 5x-25-25=5x-50=15-25).The issue is that 25 was taken from the original value of one side,but not the other,changing the value of 5x-25 by making its equality to be 15-25,then transforming 5x-25=15 to 5x-25=-10 ,wich happens whenoriginal equation's equality is unbalanced-the values change.Line 2 just applies distributive properties.